• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E.Chudakov

    Log entry time 00:01:15 on August28,2003

    Entry number 113849

    keyword=Shift summary

    Shift workers:
    E.Chudakov, J.Yuan, V.Punjabi

    16:00 kinematics 11.11 finished run 4114 (0 deg, 1/2wp out)
    16:20-17:00 0 deg,, 1/2wp in, runs 4115-4117
    17:20 Septum He level dropped and the current was automatically ramped down from 380A to 191A. It started at about 17:00.
    17:40 - we asked for control access
    17:00 - Mark Stevens and Ed Folts arrived and opened manually the valve. The He level went up.
    17:00 - NMR 41.7%
    18:20 - beam permit
    18:20-19:40 septum current ramping up
    19:40 - septum at 377.25, beam spot checked. Start run 4120
    20:50 finished 0 deg, 1/2wp in point, runs 4115-4117, 4121-4124
    21:00 start 270 deg, 1/2wp in, run 4125
    21:50 the beam spot position is unstable - jumping around every several minutes by 0.3-0.5mm, sometimes up to 1mm. MCC told they did not know why.
    22:25 finished 270 deg, 1/2wp in, runs 4125-4130
    22:27 start 270 deg,, 1/2wp out, run 4131
    22:50 the beam position became much more stable (after we complained to MCC)
    22
    23:50 finished 270 deg, 1/2wp out, runs 4131-4136
    23:50 started changing to kin.11.12 Phrs=2.685 GeV